OpenStudy (anonymous):

Find the slope and y-intercept of each line. y = x – 3 (Points : 1) slope: 3, y-intercept: –1 slope: –3, y-intercept: 1 slope: 1, y-intercept: –3 slope: –1, y-intercept: –3 help

OpenStudy (anonymous):

c i think

OpenStudy (bradely):

y = x -3 is in the form of y = mx+c Compare this m= 1 and c = -3 So slope 1 and y intercept -3 Option C is correct Source: http://www.mathskey.com/question2answer/